# RASP: A Drone-based Reconfigurable Actuation and Sensing Platform for Engaging Physical Environments with Foundation Models

### v2

@@ 5.1. Benchmarking
For the two actuation tasks, we observed an average offset of the drone, from where it was supposed to travel to drop its payload, of 9.1cm and 10.3cm for the “poison” and “medicine” tasks, respectively. The offset is on orders of centimeters, meaning the system was able to effectively deliver items to the proper location in most cases.
@@ 5.2. Reconfiguring Mid-Mission
The previous section demonstrated how RASP could be used in conjunction with static sensors in the environment to better perform tasks requiring a single sensor or actuator in a home or office setting. Here, we look at scenarios where a user issues commands that require multiple sensing modalities and actuators, highlighting how RASP could be easily reconfigured mid-mission to satisfy multi-layered tasks.
Common tasks that require RASP to reconfigure the drone mid-mission come in three different flavors: 1) actuate + actuate: multiple individual actuation tasks aggregated into one task (e.g., “bring me my medicine AND a snack”), 2) sense + sense: a task that involves multiple sensing modalities (e.g., “find me the coolest (temperature) place to sit out of the sunlight (light)”, 3) sense + actuate: tasks that involve performing actuation in response to sensing (e.g., “place rat poison (actuation) in dark areas (sensing)”). Table 4 summarizes our results, where we show the success rate of the first task (P1), the second task (P2), and the aggregate. We execute each task $20$ times. We discuss each task in more detail, next:
T2: Sense + Actuate - “Put poison in the warmest area”. In this task, the user may want to place poison (P2) for rats and bugs in warm areas (P1) where they are likely to congregate (e.g., warm places). We simulate “warm places” boiling water in a kettle in locations visible to cameras. There were three times that a location away from the kettle was chosen (P1). In these instances, our image segmentation approach cut out the area we placed the kettle, so the failure point was from the camera rather than RASP. Improving sensing with foundational models in future work is key to realizing a robust version of this end-to-end system.
5.3. In-the-Wild Deployment
@@
6. Discussion
The focus of this work is to demonstrate how a novel autonomous reconfigurable sensing and actuation drone platform is a critical step towards FMs, LLMs, and AGI that can interact, respond, and actuate our physical environments seamlessly, much like what has been achieved in the digital domain. To progress further, there are several directions that warrant further attention, which we discuss next.
Robust drone path planning, navigation, and search for indoor environments. As discussed in Section 5.3, many of the tasks that failed were due to the simple search and navigation algorithms we implemented that could not capture and detect small objects of interest in a cluttered home/indoor setting. More work into exploring, detecting, and navigating to these small objects of interest is needed.
Penetrative AI, foundation models and leveraging diverse sensors in the environment. In this work, we leveraged cameras in conjunction with language to understand and interpret the environment. However, not all areas may have a camera (e.g., due to privacy or related concerns), and there are many more types of sensors that could be deployed; enabling foundation models and drones to understand the physical world (e.g., penetrative AI (Xu et al., 2024)) through all types of sensing modalities is important for generalizing FMs and RASP to more scenarios.
@@ 7. Conclusion
We propose RASP, a modular and reconfigurable drone sensing and actuation platform that allows FMs and LLMs to dynamically to swap sensing and actuation capabilities on-the-fly. We demonstrate through a real deployment of a personal assistant application that RASP can improve the successful completion of complex tasks throughout our physical spaces by up to $85\%$. This drastic improvement stems from RASP’s ability to provide localized sensing and actuation in a large space even without large deployments of sensors and devices, as is commonly seen today. RASP is a critical step towards FMs and LLMs that can naturally interact and actuate the physical environment, just as they have shown in many applications in the digital domain.

### v3

@@ 7.2. Results and Analysis
For the two actuation tasks, we observed a median offset of the drone, from where it was supposed to travel to drop its payload, of 9.1cm and 10.3cm for the “poison” and “medicine” tasks, respectively. The offset is on orders of centimeters, meaning the system was able to effectively deliver items to the proper location in most cases.
@@ 7.3. Reconfiguring Mid-Mission
The previous section demonstrated how FlexiFly could be used in conjunction with static sensors in the environment to better perform tasks requiring a single sensor or actuator in a home or office setting. Here, we look at scenarios where a user issues commands that require multiple sensing modalities and actuators, highlighting how FlexiFly could be easily reconfigured mid-mission to satisfy multi-layered tasks.
Common tasks that require FlexiFly to reconfigure the drone mid-mission come in three different flavors: 1) actuate + actuate: multiple individual actuation tasks aggregated into one task (e.g., “bring me my medicine AND a snack”), 2) sense + sense: a task that involves multiple sensing modalities (e.g., “find me the coolest (temperature) place to sit out of the sunlight (light)”, 3) sense + actuate: tasks that involve performing actuation in response to sensing (e.g., “place rat poison (actuation) in dark areas (sensing)”). Table 4 summarizes our results, where we show the success rate of the first task (P1), the second task (P2), and the aggregate. We execute each task $20$ times. We discuss each task in more detail, next:
T2: Sense + Actuate - “Put poison in the warmest area”. In this task, the user may want to place poison (P2) for rats and bugs in warm areas (P1) where they are likely to congregate (e.g., warm places). We simulate “warm places” boiling water in a kettle in locations visible to cameras. There were three times that a location away from the kettle was chosen (P1). In these instances, our image segmentation approach cut out the area we placed the kettle, so the failure point was from the camera rather than FlexiFly. Improving sensing with foundational models in future work is key to realizing a robust version of this end-to-end system.
@@ 7.4. In-the-Wild Deployment
8. Discussion and Future Work
Usability
The deployment of autonomous drones in indoor environments raises critical usability challenges. A primary concern is noise disruption from drone propellers, which could be mitigated through: 1) flight path optimization that maintains higher operating altitudes when possible, reducing perceived ground-level noise (Adlakha et al., 2023); 2) implementation of low-noise propeller designs that recent aerodynamics research suggests could reduce noise by around 5dB (Sebastian and Strem, 2020; Jansen, 2024); and 3) context-aware navigation that avoids occupied areas during noise-sensitive periods. Beyond noise, user interaction with the system requires streamlining. We envision developing natural interaction paradigms including gesture control for intuitive drone guidance (Abtahi et al., 2017), and an augmented reality interface for visualizing drone intentions and planned paths (Mourtzis et al., 2022; Konstantoudakis et al., 2022).
Privacy
Privacy is a critical concern for camera networks and camera-equipped drones in indoor spaces. In this work, camera feeds are transmitted locally and processed on a local server for scene understanding, vision-language grounding and drone navigation. The need for server can be bypassed as more efficient, compact foundation models and powerful edge computers emerge, as well as leveraging compression techniques such as quantization and distillation to create dedicated smaller models (Sharshar et al., 2025; Cai et al., 2024) for each vision task mentioned above while preserving the generalizibility.
Extending to Diverse Environments
While we We demonstrated the effectiveness of our implementation in several indoor environments, extending FlexiFly to new settings presents unique challenges and opportunities. Flexifly can be adapted to environments with existing camera infrastructure as long as the relative position of the cameras are known; this can be achieved in a self-supervised manner by leveraging recent advances in camera self-localization and calibration techniques (Liu et al., 2022), potentially enhanced by fine-tuning vision-language models for specific deployment contexts. However, we acknowledge significant limitations in environments where camera deployment is impractical or restricted. For spaces primarily monitored by non-visual sensors (e.g., RF, vibration, or acoustic sensors), new methodologies beyond ARCK-Means and our vision pipeline must be developed for localizing areas of interest and decision-making. We envision that future FMs and penetrative AI (Xu et al., 2024) could help process diverse sensor modalities spread throughout environments, aided by novel segmentation approaches for non-visual spatial data to enable accurate drone navigation.
@@ 9. Conclusion
Our work studies the possibility of LLMs and FMs as a general intelligence for physical spaces. We identify that FMs analyzing sensing data monitoring a large space have difficulty identifying localized events that occur in small areas. As such, we propose novel segmentation methods and drones with reconfigurable sensing and actuation that enable FMs to identify and “zoom in” to analyze targeted areas with higher resolution. We demonstrate through a real deployment of a personal assistant application that FlexiFly can improve the successful completion of complex tasks throughout our physical spaces by up to $85\%$. FlexiFly is a critical step towards FMs and LLMs that can naturally interact and actuate the physical environment, just as they have shown in many applications in the digital domain.
Acknowledgements.
This research was partially supported by COGNISENSE, one of seven centers in JUMP 2.0, a Semiconductor Research Corporation (SRC) program sponsored by DARPA, as well as the National Science Foundation under Grant Number CNS-1943396. The views and conclusions contained here are those of the authors and should not be interpreted as necessarily representing the official policies or endorsements, either expressed or implied, of Columbia University, NSF, SRC, DARPA, or the U.S. Government or any of its agencies.